# Selmer stability in families of congruent Galois representations

### v1

@@ Abstract.
In this article I study the variation of Selmer groups in families of modular Galois representations that are congruent modulo a fixed prime $p\geq 5$. Motivated by analogies with Goldfeld’s conjecture on ranks in quadratic twist families of elliptic curves, I investigate the stability of Selmer groups defined over $\mathbb{Q}$ via Greenberg’s local conditions under congruences of residual Galois representations. Let $X$ be a positive real number. Fix a residual representation $\bar{\rho}$ and a corresponding modular form $f$ of weight $2$ and optimal level. I count the number of level-raising modular forms $g$ of weight $2$ that are congruent to $f$ modulo $p$, with level $N_{g}\leq X$, such that the $p$-rank of the Selmer groups of $g$ equals that of $f$. Under some mild assumptions on $\bar{\rho}$, I prove that this count grows at least as fast as $X(\log X)^{\alpha-1}$ as $X\to\infty$, for an explicit constant $\alpha>0$. The main result is a partial generalization of theorems of Ono and Skinner on rank-zero quadratic twists to the setting of modular forms and Selmer groups.
@@ 1.1. Motivation and background
The questions studied in this article are motivated by a conjecture of Dorian Goldfeld, which concerns the distribution of Mordell–Weil ranks in families of quadratic twists of a fixed elliptic curve over $\mathbb{Q}$. Given an elliptic curve $E/\mathbb{Q}$, one can consider its quadratic twists $E_{d}$, indexed by square-free integers $d$. Goldfeld conjectured that, asymptotically, 50 percent of these twists should have rank 0 and 50 percent should have rank 1, with 0 percent having rank greater than 1. This conjecture is informed by analogies with random matrix theory, the Birch and Swinnerton-Dyer conjecture, and probabilistic models for Selmer groups. An important feature of the quadratic twist family is that the mod-2 Galois representation remains fixed: for all $d$, one has an isomorphism $E_{d}[2]\simeq E[2]$ of Galois modules. This mod-2 congruence implies that the 2-Selmer groups of $E_{d}$ and $E$ are closely related—a theme that plays a central role in the work of Klagsbrun–Mazur–Rubin [KMR14], Kriz–Li [KL19], and Smith [Smi22, Smi23]. It is natural to study generalizations of this theme to $p$ congruences between modular Galois representations. I shall work with primes $p\geq 5$. Level–raising results of Diamond–Taylor [DT94] provide me with parameterizations of families of $p$-congruent Galois representations. Congruences have also played a significant role in Iwasawa theory. For a prime $p$, two elliptic curves $E$ and $E^{\prime}$ over $\mathbb{Q}$ are said to be $p$-congruent if there exists a Galois-module isomorphism $E[p]\simeq E^{\prime}[p]$. Greenberg and Vatsal [GV00] showed that for such $p$-congruent curves, the Iwasawa invariants of their $p$-primary Selmer groups over the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$ are explicitly related.
@@ 1.2. Main result
In this article, I study the behavior of Selmer groups in $p$-congruent families of modular Galois representations, with particular emphasis on their stability under congruences. I consider Selmer groups defined over $\mathbb{Q}$ by Greenberg’s local conditions and investigate analogues of phenomena observed in the context of elliptic curves, aiming to understand how congruences between Galois representations influence the global structure of Selmer groups. The Selmer groups considered are closely related to Bloch–Kato Selmer groups, however are not the same.
I fix a prime $p\geq 5$, a $p$-adic field $\mathcal{K}$ with valuation ring $\mathcal{O}$, uniformizer $\varpi$, and residue field $\kappa$. Denote by $\chi$, the $p$-adic cyclotomic character. Let $\bar{\rho}:\operatorname{G}_{\mathbb{Q}}\to\operatorname{GL}_{2}(\kappa)$ be a surjective, odd Galois representation with determinant $\bar{\chi}$, the mod-$p$ reduction of $\bar{\chi}$. Let $N_{\bar{\rho}}$ be the prime to $p$ part of the Artin conductor of $\bar{\rho}$. By results of Khare and Wintenberger [KW09a, KW09b], there exists a weight 2 Hecke eigenform $f$ of trivial nebentypus and optimal level $N_{f}=N_{\bar{\rho}}$ such that $\bar{\rho}_{f}=\bar{\rho}$. Denote by $\operatorname{Sel}_{\operatorname{Gr}}(A_{g}/\mathbb{Q})$ the Selmer group of $f$ over $\mathbb{Q}$ defined by Greenberg local conditions. This is a cofinitely generated $\varpi$-primary $\mathcal{O}$-module, which contains the $\varpi$-primary Bloch–Kato Selmer group as finite index submodule. For $X>0$, let $\mathcal{S}_{f}(X)$ denote the set of natural numbers $N\leq X$ such that $N=N_{g}$ for some Hecke eigencuspform $g$ of weight $2$ satisfying:
@@
Define $N_{f}(X):=\#\mathcal{S}_{f}(X)$.
@@
Assume that no prime ±$\ell\equiv\pm 1\pmod{p}$ divides $N_{\bar{\rho}}$. Then, with notation as above,
@@
where $\alpha:=\frac{p-3}{(p-1)^{2}}$.
@@ 2. Modular forms and their Selmer groups
In this section, I discuss modular Galois representations, congruences and associated Selmer groups. I begin by setting relevant notation.
@@ 2.1. Notation
  • Throughout, $p$ be an odd prime number and let $\chi$ be the $p$-adic cyclotomic character.
  • Let $\mathbb{F}_{p}$ be the finite field with $p$ elements. Given an $\mathbb{F}_{p}$ vector space $V$, set $V^{\vee}:=\operatorname{Hom}\left(V,\mathbb{F}_{p}\right)$.
  • Let $\operatorname{G}_{\mathbb{Q}}:=\operatorname{Gal}(\bar{\mathbb{Q}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$.
  • Let $\tau$ be a variable in the complex upper half-plane. Setting $q:=\exp(2\pi i\tau)$, let $f=\sum_{n=1}^{\infty}a_{n}(f)q^{n}$ be a normalized Hecke eigenform of weight $k\geq 2$ and level $N_{f}$. Let $F_{f}\subset\bar{\mathbb{Q}}$ denote the number field generated by its Fourier coefficients.
  • For each prime $\ell$, let $\bar{\mathbb{Q}}_{\ell}$ be a fixed algebraic closure of $\mathbb{Q}_{\ell}$, and fix an embedding $\iota_{\ell}:\bar{\mathbb{Q}}\hookrightarrow\bar{\mathbb{Q}}_{\ell}$. Let $\operatorname{G}_{\ell}:=\operatorname{Gal}(\bar{\mathbb{Q}}_{\ell}/\mathbb{Q}_{\ell})$ be the absolute Galois group of $\mathbb{Q}_{\ell}$. I adopt the short hand notation $H^{i}(\mathbb{Q}_{\ell},\cdot):=H^{i}(\operatorname{G}_{\ell},\cdot)$. The embedding $\iota_{\ell}$ induces an inclusion
    $\iota_{\ell}^{*}:\operatorname{G}_{\ell}\hookrightarrow\operatorname{G}_{\mathbb{Q}}.$
  • The inertia subgroup of $\operatorname{G}_{\ell}$ is denoted by $\operatorname{I}_{\ell}$ and let $\sigma_{\ell}\in\operatorname{G}_{\ell}/\operatorname{I}_{\ell}$ be the Frobenius element. When there is no cause for confusion, I occasionally choose a lift in $\operatorname{G}_{\ell}$ of the Frobenius, which I shall also denote by $\sigma_{\ell}$.
@@
    $H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},M):=\operatorname{image}\left\{H^{1}(\operatorname{G}_{\ell}/\operatorname{I}_{\ell},M^{\operatorname{I}_{\ell}})\longrightarrow H^{1}(\mathbb{Q}_{\ell},M)\right\}.$
@@
  • The embedding $\iota_{\ell}$ induces an injection $\iota_{\ell}^{*}:\operatorname{G}_{\ell}\hookrightarrow\operatorname{G}_{\mathbb{Q}}$. Let $\mathcal{K}/\mathbb{Q}_{p}$ be the finite extension generated by $\iota_{p}(F_{f})$, let $\mathcal{O}$ be the valuation ring of $\mathcal{K}$ and $\varpi$ be a uniformizer of $\mathcal{O}$. Set $\kappa$ to denote the residue field $\mathcal{O}/(\varpi)$. If $\rho$ is a representation of $\operatorname{G}_{\mathbb{Q}}$, denote by $\rho_{|\ell}$ the restriction of $\rho$ to $\operatorname{G}_{p}$. Then $\rho$ is unramified at $\ell$ if $\rho_{|\operatorname{I}_{\ell}}$ is trivial. In this case, $\rho(\sigma_{\ell})$ is well defined.
@@
  • Given a finite set of primes $S$, let $\mathbb{Q}_{S}\subset\bar{\mathbb{Q}}$ be the maximal extension of $\mathbb{Q}$ in which all primes $\ell\notin S$ are unramified. Set $\operatorname{G}_{S}:=\operatorname{Gal}(\mathbb{Q}_{S}/\mathbb{Q})$ and $H^{i}(\mathbb{Q}_{S}/\mathbb{Q},\cdot):=H^{i}(\operatorname{G}_{S},\cdot)$.
2.2. Facts from Galois cohomology
@@
Let $S$ be a finite set of rational primes which includes $p$ and suppose $M$ is a finite-dimensional vector space over $\mathbb{F}_{p}$, equipped with an action of $\operatorname{G}_{S}$. For $i=1,2$, define
@@
(${}_{S}^{i}(M):=\ker\left(H^{i}(\mathbb{Q}_{S}/\mathbb{Q},M)\longrightarrow\bigoplus_{\ell\in S}H^{i}(\mathbb{Q}_{\ell},M)\right).$
@@
Now, for each prime $\ell\in S$, fix a choice of subspace $\mathcal{L}_{\ell}\subseteq H^{1}(\mathbb{Q}_{\ell},M)$. The collection ={}$\mathcal{L}=\{\mathcal{L}_{\ell}\}_{\ell\in S}$ specifies a local condition at each prime. Using these local conditions, one defines the Selmer group $H^{1}_{\mathcal{L}}(\mathbb{Q}_{S}/\mathbb{Q},M)$ as follows:
@@
$H^{1}_{\mathcal{L}}(\mathbb{Q}_{S}/\mathbb{Q},M):=\ker\left(H^{1}(\mathbb{Q}_{S}/\mathbb{Q},M)\longrightarrow\bigoplus_{\ell\in S}\frac{H^{1}(\mathbb{Q}_{\ell},M)}{\mathcal{L}_{\ell}}\right).$
@@
One introduces the dual module $M^{*}:=\operatorname{Hom}(M,\mu_{p})$, where $\mu_{p}$ denotes the group of $p$-th roots of unity (viewed as a Galois module). Poitou–Tate global duality, provides an identification
(2.1) (${}_{S}^{2}(M)\simeq{}_{S}^{1}(M^{*})^{\vee}.$
At the local level, for each prime $\ell\in S$, the local Tate pairing provides a perfect, nondegenerate pairing
$H^{1}(\mathbb{Q}_{\ell},M)\times H^{1}(\mathbb{Q}_{\ell},M^{*})\longrightarrow H^{2}(\mathbb{Q}_{\ell},\mu_{p})\simeq\mathbb{F}_{p}.$
Using this pairing, for each $\ell\in S$, the orthogonal complement $\mathcal{L}_{\ell}^{\perp}\subseteq H^{1}(\mathbb{Q}_{\ell},M^{*})$ of $\mathcal{L}_{\ell}$ is defined to consist of all elements of $H^{1}(\mathbb{Q}_{\ell},M^{*})$ that pair trivially with every element of $\mathcal{L}_{\ell}$.
@@
Correspondingly, the dual Selmer group $H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})$ is defined to be the subgroup of $H^{1}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})$ consisting of classes whose local restrictions lie inside the subspaces $\mathcal{L}_{\ell}^{\perp}$ for each $\ell\in S$, that is,
@@
(2.2) $H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},M^{*}):=\ker\left(H^{1}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})\longrightarrow\bigoplus_{\ell\in S}\frac{H^{1}(\mathbb{Q}_{\ell},M^{*})}{\mathcal{L}_{\ell}^{\perp}}\right).$
Wiles’ formula (see [NSW08, Theorem 8.7.9]), relates the dimensions of a Selmer group and its dual. Explicitly,
$\begin{split}\dim_{\mathbb{F}_{p}}&H^{1}_{\mathcal{L}}(\mathbb{Q}_{S}/\mathbb{Q},M)-\dim_{\mathbb{F}_{p}}H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})\\ =&\dim_{\mathbb{F}_{p}}H^{0}(\mathbb{Q},M)-\dim_{\mathbb{F}_{p}}H^{0}(\mathbb{Q},M^{*})+\sum_{\ell\in S\cup\{\infty\}}\left(\dim_{\mathbb{F}_{p}}\mathcal{L}_{\ell}-\dim_{\mathbb{F}_{p}}H^{0}(\mathbb{Q}_{\ell},M)\right),\\ \end{split}$
where it is understood that $\mathcal{L}_{\infty}=0$. The Poitou–Tate exact sequence gives the following:
(2.3) $0\longrightarrow H^{1}_{\mathcal{L}}(\mathbb{Q}_{S}/\mathbb{Q},M)\longrightarrow H^{1}(\mathbb{Q}_{S}/\mathbb{Q},M)\longrightarrow\bigoplus_{\ell\in S}\left(\frac{H^{1}(\mathbb{Q}_{\ell},M)}{\mathcal{L}_{\ell}}\right)\longrightarrow H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})^{\vee}\longrightarrow{}_{S}^{2}(M)\longrightarrow 0,$
see, for example, [Tay03, p. 555, line 7].

### v2

@@ Abstract.
In this article, we study the variation of Selmer groups in families of modular Galois representations that are congruent modulo a fixed prime $p\geq 5$. Motivated by analogies with Goldfeld’s conjecture on ranks in quadratic twist families of elliptic curves, we investigate the stability of Selmer groups defined over $\mathbb{Q}$ via Greenberg’s local conditions under congruences of residual Galois representations. Let $X$ be a positive real number. Fix a residual representation $\bar{\rho}$ and a corresponding modular form $f$ of weight $2$ and optimal level. We count the number of level-raising modular forms $g$ of weight $2$ that are congruent to $f$ modulo $p$, with level $N_{g}\leq X$, such that the $p$-rank of the Selmer groups of $g$ equals that of $f$. Under some mild assumptions on $\bar{\rho}$, we prove that this count grows at least as fast as $X(\log X)^{\alpha-1}$ as $X\to\infty$, for an explicit constant $\alpha>0$. The main result is a partial generalization of theorems of Ono and Skinner on rank-zero quadratic twists to the setting of modular forms and Selmer groups.
@@ 1.1. Motivation and background
The questions studied in this article are motivated by a conjecture of Goldfeld [GOL79], which concerns the distribution of Mordell–Weil ranks in families of quadratic twists of a fixed elliptic curve over $\mathbb{Q}$. Given an elliptic curve $E/\mathbb{Q}$, one can consider its quadratic twists $E_{d}$, indexed by square-free integers $d$. Goldfeld conjectured that, asymptotically, 50 percent of these twists should have rank 0 and 50 percent should have rank 1, with 0 percent having rank greater than 1. This conjecture is informed by analogies with random matrix theory, the Birch and Swinnerton-Dyer conjecture, and probabilistic models for Selmer groups. An important feature of the quadratic twist family is that the mod-2 Galois representation remains fixed: for all $d$, one has an isomorphism $E_{d}[2]\simeq E[2]$ of Galois modules. This mod-2 congruence implies that the 2-Selmer groups of $E_{d}$ and $E$ are closely related—a theme that plays a central role in the work of Klagsbrun–Mazur–Rubin [KMR14], Kriz–Li [KL19], and Smith [SMI22, SMI23]. It is natural to study generalizations of this theme to $p$ congruences between modular Galois representations. We shall work with primes $p\geq 5$. Level–raising results of Diamond–Taylor [DT94] provide us with parameterizations of families of $p$-congruent Galois representations. Congruences have also played a significant role in Iwasawa theory. For a prime $p$, two elliptic curves $E$ and $E^{\prime}$ over $\mathbb{Q}$ are said to be $p$-congruent if there exists a Galois-module isomorphism $E[p]\simeq E^{\prime}[p]$. Greenberg and Vatsal [GV00] showed that for such $p$-congruent curves, the Iwasawa invariants of their $p$-primary Selmer groups over the cyclotomic $\mathbb{Z}_{p}$-extension of $\mathbb{Q}$ are explicitly related.
@@ 1.2. Main result
In this article, we study the behavior of Selmer groups in $p$-congruent families of modular Galois representations, with particular emphasis on their stability under congruences. We consider Selmer groups defined over $\mathbb{Q}$ by Greenberg’s local conditions and investigate analogues of phenomena observed in the context of elliptic curves, aiming to understand how congruences between Galois representations influence the global structure of Selmer groups. The Selmer groups considered are closely related to Bloch–Kato Selmer groups, however are not the same.
We fix a prime $p\geq 5$, a $p$-adic field $\mathcal{K}$ with valuation ring $\mathcal{O}$, uniformizer $\varpi$, and residue field $\kappa$. Denote by $\chi$, the $p$-adic cyclotomic character. Let $\bar{\rho}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\to\operatorname{GL}_{2}(\kappa)$ be a surjective, odd Galois representation with determinant $\bar{\chi}$, the mod-$p$ reduction of $\chi$. Let $N_{\bar{\rho}}$ be the prime to $p$ part of the Artin conductor of $\bar{\rho}$. By results of Khare and Wintenberger [KW09a, KW09b], there exists a weight 2 Hecke eigenform $f$ of trivial nebentypus and optimal level $N_{f}=N_{\bar{\rho}}$ such that $\bar{\rho}_{f}=\bar{\rho}$. Associated with $f$ is a modular Galois representation $\rho_{f}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\operatorname{Aut}_{\mathcal{K}}(V_{f})\xrightarrow{\sim}\operatorname{GL}_{2}(\mathcal{K})$, where $V_{f}\simeq\mathcal{K}^{2}$ is the underlying vector space. There exists a Galois stable lattice $T_{f}\subset V_{f}$. We set $A_{f}\mathrel{\mathop{\ordinarycolon}}=V_{f}/T_{f}\simeq(\mathcal{K}/\mathcal{O})^{2}$. Denote by $\operatorname{Sel}_{\operatorname{Gr}}(A_{f}/\mathbb{Q})$ the Selmer group of $f$ over $\mathbb{Q}$ defined by Greenberg local conditions. This is a cofinitely generated $\varpi$-primary $\mathcal{O}$-module, which contains the $\varpi$-primary Bloch–Kato Selmer group as a finite index submodule. For $X>0$, let $\mathcal{S}_{f}(X)$ denote the set of natural numbers $N\leq X$ such that $N=N_{g}$ for some Hecke eigencuspform $g$ of weight $2$ satisfying:
@@
Define $N_{f}(X)\mathrel{\mathop{\ordinarycolon}}=\#\mathcal{S}_{f}(X)$.
@@
With respect to notation as above, assume that no prime $\ell\equiv\pm 1\pmod{p}$ divides $N_{\bar{\rho}}$. Then,
@@
where $\alpha\mathrel{\mathop{\ordinarycolon}}=\frac{p-3}{(p-1)^{2}}$.
@@
For $2$-Selmer groups of elliptic curves in quadratic twist families of elliptic curves over number fields, results similar to the Theorem above have been proven by Mazur and Rubin (see [MR10, Theorem 1.4]).
@@
Acknowledgement
We thank the anonymous referee for the excellent report.
@@ 2. Modular forms and their Selmer groups
In this section, we discuss modular Galois representations, congruences and associated Selmer groups. We begin by setting relevant notation.
@@ 2.1. Notation
  • Throughout, $p$ will be an odd prime number.
  • Let $\mathbb{F}_{p}$ be the finite field with $p$ elements. Given an $\mathbb{F}_{p}$ vector space $V$, set $V^{\vee}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Hom}\left(V,\mathbb{F}_{p}\right)$.
  • Let $\operatorname{G}_{\mathbb{Q}}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\bar{\mathbb{Q}}/\mathbb{Q})$ be the absolute Galois group of $\mathbb{Q}$.
  • Denote by $\chi\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\mathbb{Q}}\rightarrow\mathbb{Z}_{p}^{\times}$ the $p$-adic cyclotomic character and let $\bar{\chi}$ be its mod-$p$ reduction.
  • Let $\tau$ be a variable in the complex upper half-plane. Setting $q\mathrel{\mathop{\ordinarycolon}}=\exp(2\pi i\tau)$, let $f=\sum_{n=1}^{\infty}a_{n}(f)q^{n}$ be a normalized Hecke eigenform of weight $k\geq 2$ and level $N_{f}$. Let $F_{f}\subset\bar{\mathbb{Q}}$ denote the number field generated by its Fourier coefficients.
  • For each prime $\ell$, let $\bar{\mathbb{Q}}_{\ell}$ be a fixed algebraic closure of $\mathbb{Q}_{\ell}$, and fix an embedding $\iota_{\ell}\mathrel{\mathop{\ordinarycolon}}\bar{\mathbb{Q}}\hookrightarrow\bar{\mathbb{Q}}_{\ell}$. Let $\operatorname{G}_{\ell}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\bar{\mathbb{Q}}_{\ell}/\mathbb{Q}_{\ell})$ be the absolute Galois group of $\mathbb{Q}_{\ell}$. We adopt the short hand notation $H^{i}(\mathbb{Q}_{\ell},\cdot)\mathrel{\mathop{\ordinarycolon}}=H^{i}(\operatorname{G}_{\ell},\cdot)$. The embedding $\iota_{\ell}$ induces an inclusion
    $\iota_{\ell}^{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\ell}\hookrightarrow\operatorname{G}_{\mathbb{Q}}.$
  • The inertia subgroup of $\operatorname{G}_{\ell}$ is denoted by $\operatorname{I}_{\ell}$ and let $\sigma_{\ell}\in\operatorname{G}_{\ell}/\operatorname{I}_{\ell}$ be the Frobenius element. When there is no cause for confusion, we occasionally choose a lift in $\operatorname{G}_{\ell}$ of the Frobenius, which we shall also denote by $\sigma_{\ell}$.
@@
    $H^{1}_{\operatorname{nr}}(\mathbb{Q}_{\ell},M)\mathrel{\mathop{\ordinarycolon}}=\operatorname{image}\left\{H^{1}(\operatorname{G}_{\ell}/\operatorname{I}_{\ell},M^{\operatorname{I}_{\ell}})\longrightarrow H^{1}(\mathbb{Q}_{\ell},M)\right\}.$
@@
  • The embedding $\iota_{\ell}$ induces an injection $\iota_{\ell}^{*}\mathrel{\mathop{\ordinarycolon}}\operatorname{G}_{\ell}\hookrightarrow\operatorname{G}_{\mathbb{Q}}$. Let $\mathcal{K}/\mathbb{Q}_{p}$ be the finite extension generated by $\iota_{p}(F_{f})$, let $\mathcal{O}$ be the valuation ring of $\mathcal{K}$ and $\varpi$ be a uniformizer of $\mathcal{O}$. Set $\kappa$ to denote the residue field $\mathcal{O}/(\varpi)$. If $\rho$ is a representation of $\operatorname{G}_{\mathbb{Q}}$, denote by $\rho_{|\ell}$ the restriction of $\rho$ to $\operatorname{G}_{p}$. Then $\rho$ is unramified at $\ell$ if $\rho_{|\operatorname{I}_{\ell}}$ is trivial. In this case, $\rho(\sigma_{\ell})$ is well defined.
@@
  • Given a finite set of primes $S$, let $\mathbb{Q}_{S}\subset\bar{\mathbb{Q}}$ be the maximal extension of $\mathbb{Q}$ in which all primes $\ell\notin S$ are unramified. Set $\operatorname{G}_{S}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Gal}(\mathbb{Q}_{S}/\mathbb{Q})$ and $H^{i}(\mathbb{Q}_{S}/\mathbb{Q},\cdot)\mathrel{\mathop{\ordinarycolon}}=H^{i}(\operatorname{G}_{S},\cdot)$.
2.2. Selmer and dual Selmer data
@@
Let $S$ be a finite set of rational primes which includes $p$ and suppose $M$ is a finite-dimensional vector space over $\mathbb{F}_{p}$, equipped with a continuous action of $\operatorname{G}_{S}$. For $i=1,2$, define
@@
$_{S}^{i}(M)\mathrel{\mathop{\ordinarycolon}}=\ker\left(H^{i}(\mathbb{Q}_{S}/\mathbb{Q},M)\longrightarrow\bigoplus_{\ell\in S}H^{i}(\mathbb{Q}_{\ell},M)\right).$
@@
Now, for each prime $\ell\in S$, fix a choice of subspace $\mathcal{L}_{\ell}\subseteq H^{1}(\mathbb{Q}_{\ell},M)$. The collection $\mathcal{L}=\{\mathcal{L}_{\ell}\}_{\ell\in S}$ specifies a local condition at each prime. Using these local conditions, one defines the Selmer group $H^{1}_{\mathcal{L}}(\mathbb{Q}_{S}/\mathbb{Q},M)$ as follows:
@@
$H^{1}_{\mathcal{L}}(\mathbb{Q}_{S}/\mathbb{Q},M)\mathrel{\mathop{\ordinarycolon}}=\ker\left(H^{1}(\mathbb{Q}_{S}/\mathbb{Q},M)\longrightarrow\bigoplus_{\ell\in S}\frac{H^{1}(\mathbb{Q}_{\ell},M)}{\mathcal{L}_{\ell}}\right).$
@@
One introduces the dual module $M^{*}\mathrel{\mathop{\ordinarycolon}}=\operatorname{Hom}(M,\mu_{p})$, where $\mu_{p}$ denotes the group of $p$-th roots of unity (viewed as a Galois module).
@@
With respect to the Tate pairing, for each $\ell\in S$, the orthogonal complement $\mathcal{L}_{\ell}^{\perp}\subseteq H^{1}(\mathbb{Q}_{\ell},M^{*})$ of $\mathcal{L}_{\ell}$ is defined to consist of all elements of $H^{1}(\mathbb{Q}_{\ell},M^{*})$ that pair trivially with every element of $\mathcal{L}_{\ell}$. The dual Selmer group $H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})$ is defined to be the subgroup of $H^{1}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})$ consisting of classes whose local restrictions lie inside the subspaces $\mathcal{L}_{\ell}^{\perp}$ for each $\ell\in S$, that is,
@@
(2.1) $H^{1}_{\mathcal{L}^{\perp}}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})\mathrel{\mathop{\ordinarycolon}}=\ker\left(H^{1}(\mathbb{Q}_{S}/\mathbb{Q},M^{*})\longrightarrow\bigoplus_{\ell\in S}\frac{H^{1}(\mathbb{Q}_{\ell},M^{*})}{\mathcal{L}_{\ell}^{\perp}}\right).$